Operations Report – January 1st

Crew 218 Operations Report 1-Jan-20

SOL: 10

Name of person filing report: Luz Maria Agudelo Urrego

Non-nominal systems: NA

Generator: run

Hours run: 12

From what time last night: 1930

To what time this morning: 0730

List any additional daytime hours when the generator was run: N/A

Solar— SOC % (Before the generator is run at night: 62%

Diesel Reading – 74%

Station Propane Reading – 65%

Ethanol Free Gasoline: N/A

Water (loft tank) (gal): 55

Water Meter (units): 0147212.1

Water (static tank) (gal): 270

Static to Loft Pump used – No

Water in Green Hab (gal): 84

Water in Science Dome (gal): 0

Toilet tank emptied: Yes

Deimos rover used: Still in the workshop

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: Assigned to director

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: EVA 8

Hours: 121.5

Beginning charge: (Before EVA): 100%

Ending charge: (On return from EVA, before recharging): 29%

Currently charging: Yes

Spirit rover used: EVA 9 – Rover was not used

Hours: 121.5

Beginning charge: (Before EVA):

Ending charge: (On return from EVA, before recharging):

Currently charging: Yes

Opportunity rover used: Still in the workshop

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: EVA 8

Hours: 126.9

Beginning charge: 100%

Ending charge: (On return from EVA, before charging): 40%

Currently charging: Yes

Curiosity rover used: EVA 9 – Rover was not used

Hours: 126.9

Beginning charge:

Ending charge: (On return from EVA, before charging):

Currently charging: Yes

Notes on rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use: N/A

Oil Added? No

# Hours the ATVs were used today: 0

Notes on ATVs: N/A

HabCar used, and why, where? No

CrewCar used, and why, where? No

General notes and comments: The tablespoon in the Hab kitchen is broken.

Summary of the internet: Nominal

Summary of suits and radios: On the two-piece space simulator Suit N°6, the right air pipe connector sheared.

Summary of Hab operations: The pipe from the static tank through the loft tank is still frozen. The crew filled out the Loft Tank by carrying water buckets from the Static Tank again.

Summary of Science Dome operations: Nominal

Summary of RAM operations: Nominal

Summary of any observatory issues: The fire alarm went on at 1820. The crew replaced the batteries in it.

Summary of health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: NA